What is the electric field due to a point charge of 20µC at a distance of 1 meter away from it?

How much charge experiences a force of 5.0 x10-3 N due to an electric field of 10 N/C?


1
Expert's answer
2022-03-08T17:02:34-0500

Electric field

"E=\frac{kq}{r^²}"


"E=\frac{9\times10^9\times20\times10^{-6}}{1^2}=18\times10^4N/C"

Force

"F=qE"

"q=\frac{F}{E}"

"q=\frac{5\times10^{-3}}{10}=5\times10^{-4}C"